A traveling wave on a taut string with a tension force T is given by the wave function: y(x,t) = 0.1sin⁡(2πx-300t), where x and y are in meters and t is in seconds. The linear mass density of the string is μ = 10 g/m, and the string is 10 m-long. The total energy on the string is



1
Expert's answer
2021-06-17T15:08:00-0400

The total energy on the string is


"E=\frac12\mu\omega^2A^2\lambda."

As we see from the equation,

"\omega=300\text{ s}^{-1},\\
A=0.1\text{ m},\\
\lambda=\frac{2\pi}{k}=\frac{2\pi}{2\pi}=1\text{ m}."


Therefore:


"E=\frac12·0.01·(300)^2(0.1)^2(1)=4.5\text{ J}."